One of the residuals in a linear regression model is equal to 6.5. Other results from the model are MSE = 3.2, MSR = 4.5, SSE = 182, and SSR = 243. What is the value of the standardized residual?
Answer #1

df error =SSE/MSE = 182/3.2 = 57

df regression = SSR/MSR = 243/4.5 = 54standardized residual